Two angles are complementary. One angle is 12degrees larger than the other . Using two variables x and y, find the size of each angle by solving a system of equations.
When I was trying to figure this out I got the answer x=51 and y=39
I don't think that is write....
Can you show me how to this equation should be wrote out please and thank you
Jeanine

Answer by stanbon(75887)

Two angles are complementary. One angle is 12degrees larger than the other . Using two variables x and y, find the size of each angle by solving a system of equations.
-------
Equations::
x + y = 90 degrees
x = y + 12
-------
Substitute for "x" and solve for "y"::
y+12 + y = 90
2y = 78
y = 39 degrees
----
x = 90-y = 90-39 = 51 degrees
------------
You are correct.
Cheers,
Stan H.